OpenStudy (anonymous):

Tyler has two fish tanks in the shape of rectangular prisms. Each of the sides of the second fish tank is 6 times greater than the sides of the first fish tank. If Tyler fills both the fish tanks completely with water, the volume of water in the second fish tank will be 216 times the volume of water in the first fish tank 6 times the volume of water in the first fish tank 36 times the volume of water in the first fish tank 12 times the volume of water in the first fish tank

OpenStudy (hoblos):

6^3 = 216

OpenStudy (hoblos):

V = length*width*height l' = 6l w'=6w h'=6h V'= l*w*h = 6l*6w*6h = 216*l*w*h = 216V